# Hierarchical Safety Realignment: Lightweight Restoration of Safety in Pruned Large Vision-Language Models

### v1

@@ Abstract
With the increasing size of Large Vision-Language Models (LVLMs), network pruning techniques aimed at compressing models for deployment in resource-constrained environments have garnered significant attention. However, we observe that pruning often leads to a degradation in safety performance. To address this issue, we present a novel and lightweight approach, termed Hierarchical Safety Realignment (HSR). HSR operates by first quantifying the contribution of each attention head to safety, identifying the most critical ones, and then selectively restoring neurons directly within these attention heads that play a pivotal role in maintaining safety. This process hierarchically realigns the safety of pruned LVLMs, progressing from the attention head level to the neuron level. We validate HSR across various models and pruning strategies, consistently achieving notable improvements in safety performance. To our knowledge, this is the first work explicitly focused on restoring safety in LVLMs post-pruning. The code will be available at https://github.com/TheShineyue/HSR.
Hierarchical Safety Realignment: Lightweight Restoration of Safety in Pruned Large Vision-Language Models
Yue Lithanks:   Equal contribution.1, Xin Yi11footnotemark: 11, Dongsheng Shi1, Gerard de Melo2, Xiaoling Wang1, Linlin Wangthanks:   Corresponding author.1 1East China Normal University, China 2University of Potsdam, Germany {yue_li, xinyi, dongsheng}@stu.ecnu.edu.cn, demelo@uni-potsdam.de, {xlwang, llwang}@cs.ecnu.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) benefit from their large parameter size and excellent model architecture, and have achieved outstanding results on diverse benchmarks. Building on this success, efforts to extend LLMs into multimodal domains have as well witnessed remarkable progress. Currently, most Large Vision-Language Models (LVLMs) consist primarily of visual encoders, adapters, and LLM backbones (Liu et al., 2024a). These models leverage image-text datasets to achieve effective multimodal alignment. To enable model deployment and application under resource-constrained environments, pruning methods (Sun et al., 2024; Frantar and Alistarh, 2023; Lee et al., 2018) calculate the importance scores of neurons to eliminate neurons that are considered rather unimportant, thereby reducing the size of the model while retaining the utility of the model to the extent possible. This line of work has seen widespread adoption for improving model efficiency.
@@
Figure 1: The Average ASR (lower values denote higher safety) of full baselines versus pruned models (50% sparsity) across safety evaluation benchmarks. The abbreviations on the x-axis, arranged from left to right, correspond to the following models: Qwen2-VL, Qwen2.5-VL, LLaVA-NeXT-Mistral, LLaVA-NeXT-Vicuna, LLaVA-NeXT-Llama3, and Llama3.2-Vision.
@@
Safety is often defined as preventing models from following malicious instructions and generating toxic content (Bianchi et al., 2023). Recent work (Zhou et al., 2025) has found that certain safety heads within the attention mechanism are crucial for feature integration in safety tasks. Additionally, neural-level research (Wei et al., 2024) has found that certain regions in the model are critical for safety guardrails, which are separate from the utility-related regions and are sparse. A natural concern arises from the fact that, because these regions contribute minimally to utility, they are prone to being removed by pruning technologies that prioritize utility importance as a pruning metric. This removal could result in a decline in the safety of the pruned model. To verify whether this problem exists, we used the Wanda pruning method to prune six mainstream LVLMs and compared their safety changes before and after pruning. The experimental results shown in Figure 1 reveal that all LVLMs exhibited varying degrees of safety degradation, with the worst-performing model showing a 15.4% safety drop and the best-performing case exhibiting a 2.7% decline. Despite the serious safety risks of pruning technologies, research on model safety restoration after pruning remains scarce.
@@
To address this problem, this paper proposes the Hierarchical Safety Realignment (HSR) approach, designed to restore the safety performance degraded by pruning, without significantly increasing the model’s parameter size. HSR hierarchically realigns the safety of the pruned model from the attention head level to the neuron level. This approach operates in two steps: First, we evaluate each attention head’s contribution to model safety and identify safety-critical heads with the greatest impact. Subsequently, for these key heads, we pinpoint and restore safety-critical neurons that should have been preserved during pruning, effectively realigning the model’s safety alignment.
@@
We have validated the effectiveness of our approach across various models and pruning techniques. Our proposed HSR approach successfully realigns the safety of pruned models, restoring over 27% of the lost safety in many cases and more than 14% even in the worst-case scenarios, all with lightweight modifications. Furthermore, through extensive analysis and ablation experiments, we have uncovered several key insights into model safety. These include the finding that a small subset of neurons plays a disproportionately significant role in ensuring safety, and the observation that certain neurons exist which negatively impact safety.
@@
  • We propose a novel method for realigning the safety of pruned LVLMs, called Hierarchical Safety Realignment (HSR). This approach achieves substantial safety improvements with lightweight modifications. To the best of our knowledge, HSR is the first method specifically designed to address the safety realignment of pruned LVLMs.
  • We validate the proposed method across various LVLMs with different pruning techniques. Extensive experiments demonstrate the superiority of our approach, showing consistent performance improvements with minimal neuron restoration.
@@ 2.2.1 Quantifying Neuron Importance
    This equation is the first-order Taylor approximation to the change of the loss when the weight entry $\mathbf{W}_{ij}$ is set to zero. Following the experimental setup described by Wei et al. (2024), for a given calibration dataset $D$, we have:
@@ 3 Experimental Setup
During the evaluation phase, we employ the following benchmarks: (1) For utility assessment, MMbench (Liu et al., 2025) and DocVQA (Mathew et al., 2021) are utilized; (2) For safety evaluation, safebench-mini (Xu et al., 2022) and the harmful subset of $\text{C}h^{3}\text{Ef}$ (Shi et al., 2024) are adopted. To ensure fair comparison and reproducibility, all evaluations are conducted under strict zero-shot settings with greedy decoding strategies.
@@
Our experiments involved six mainstream LVLMs, including three variants of LLaVA-NeXT (Li et al., 2024), built on different language models: Vicuna, Mistral, and Llama3. Additionally, we evaluated Qwen2.5-VL (Team, 2025), Qwen2-VL (Wang et al., 2024a), and Llama-3.2-Vision (Dubey et al., 2024). All models have parameter counts ranging from 7B to 11B.
@@
Three evaluation metrics we use are Attack Success Rate (ASR), Average Normalized Levenshtein Similarity (ANLS) and Accuracy (Acc). ASR is used to evaluate the safety of the model. The smaller the ASR, the better the safety. We use Llama-Guard-3-Vision (Chi et al., 2024) to determine whether the response is safe. We use ANLS and Acc as the evaluation indicators of DocVQA and MMbench respectively. Following Mathew et al.’s (2021) proposal, ANLS can ensure that minor answer mismatches stemming from OCR errors are not severely penalized. Additionally, in order to more intuitively and fairly reflect the performance improvement of our method, we provide the Ratio of model Safety Realignment (RSR) as follows:
@@
to quantify the ratio of restored safety to lost safety.
We report more details in Appendix  A.
@@ 4 Experimental Results
Regarding utility, both Qwen2-VL and the aforementioned Qwen2.5-VL show improved performance, while other models experience slight declines. This difference may be due to the Qwen series’ grouped query attention, which features the largest number of queries per group and the fewest heads (as shown in Appendix B), making each neuron’s contribution (both utility and safety) particularly significant.
Regarding the neuron restoration ratio, for LLaVA-NeXT-Vicuna, it is significantly higher than for others. This may be because other models are based on the group attention mechanism, and the benefits brought by the modification can affect a wider range.
@@
Moreover, the utility of realigned models exhibits a slight reduction at 40% sparsity while showing progressive improvement at 50 - 60% sparsity levels, peaking at 60% sparsity. This phenomenon may stem from safety-critical neurons possessing greater safety significance than utility value, whereas in heavily pruned models (50 - 60% sparsity) with substantial utility degradation, the utility of the safety-critical neurons can still bring about some improvement.
@@
Next, we analyze the impact of the hyper-parameter $\text{p}_{\max}$ as shown in Table 6, The results reveal that utility initially increases and then decreases, reaching its peak at 0.55. This underscores the importance of carefully designing $\text{p}_{\max}$ to regulate safety-related neurons. Regarding the trend of safety first deteriorating and then improving, this suggests that some neurons contribute significantly to both safety and utility, while others contribute minimally to both. By selecting an appropriate $\text{p}_{\max}$, we can effectively exclude the latter group.
@@
Finally, we examine the impact of the hyper-parameter $h$ as presented in Table 7. Since each group attention head in LLaVA-NeXT-Mistral corresponds to 4 query matrices, experiments are conducted in multiples of 4. The results show that safety performs best at $\text{h}=4$, followed by significant fluctuations. Utility also shows considerable variability. This may be due to the uneven distribution of neurons across heads: some contribute strongly to both safety and utility, while others contribute minimally. This highlights the need for more fine-grained control to address this variability effectively.
@@
Lightweight Implementation HSR can achieve safety realignment of pruned model while preserving the current sparsity of the pruned model as much as possible. For Qwen2.5-VL and LLaVA-NeXT-Llama3, the scale of neurons restored by HSR-a is 18 and 650 times that of HSR, respectively.
Enhanced Effectiveness Certain neurons have been found to adversely affect safety. By filtering them out through set operations, we ensure robust safety realignment efficacy. Directly repairing the entire head makes the safety worse (ASR increased > 0.2). This also shows again that there may be some neurons that are harmful to safety.
@@ 5 Further Analysis
We analyzed the total Ships of each model and the extent of safety degradation after pruning, finding a strong positive correlation between them. Spearman’s rank correlation coefficient $\rho$ (Spearman, 1961) is a nonparametric statistical test measuring the correlation between the ranks of two variables:
@@ 6 Related Work
Many studies have researched methods to compromise the safety of LVLMs. Gong et al. (2023) introduced FigStep, which converts prohibited content into images using typesetting to bypass safety alignment. Ying et al. (2024) proposed BAP, a jailbreak attack method that jointly optimizes text and visual prompts. In response, research on defending against such attacks and enhancing model safety has also emerged. Liu et al. (2024b) improved defense against harmful images by incorporating a security module via a two-stage training process. Meanwhile, Wang et al. (2024b) proposed AdaShield, which protects Multimodal Large Language Models from structure-based jailbreak attacks by adding a defense hint to the input, without requiring model fine-tuning or additional module training.
@@
As model sizes continue to grow, pruning techniques (Sung et al., 2024; Jianjian et al., 2024) for compressing neural networks by removing neurons have attracted significant attention. These techniques can be broadly categorized into structured and unstructured pruning. Structured pruning (Ashkboos et al., 2024; An et al., 2024) has the advantage of accelerating pruned models on standard hardware without relying on specialized support (Zhu et al., 2024), while unstructured pruning (Lee et al., 2019; Sun et al., 2024; Frantar and Alistarh, 2023) helps preserve performance at higher sparsity levels. For pruned models, Jin et al. (2022) observed that pruning introduces additional regularization, reducing accuracy loss on noisy examples in dense models. Hasan et al. (2024) noted improved model safety at low sparsity, attributing it to sharper attention. However, the safety degradation caused by pruning at slightly higher sparsity has been overlooked, motivating our research on methods to realign the safety of pruned models.
@@ 7 Conclusion
This paper studies the problem of safety degradation after pruning LVLMs, and proposes a new HSR method, which achieves the goal of realigning the safety of the pruned model by first identifying the safety-critical attention heads that contribute greatly to safety at the attention head level, and then identifying and restoring the safety-critical neurons that should have been pruned on the safety-critical attention heads. To our knowledge, HSR is the first study to focus on the problem of pruned model restoration. Our HSR achieves effective safety restoration for a range of mainstream LVLMs and pruning methods, and only requires a very small scale of neuron restoration. We further provide an explanation of the internal structure of the model through experiments. We hope that the HSR method can help deploy safe and compact LVLMs.
@@ Appendix E Analysis and visualization of the overlap between $S^{s}(\text{q})$ and $S^{u}(\text{p})$
  • For Q, K, V and O matrices, the degree of differentiation between safety and utility increases initially and then decreases as the layers deepen. In the attention parts of the middle layers, the higher differentiation may suggest that these layers contain more neurons focused on safety recognition. This observation is similar to the findings from prior work (Arditi et al., ), where it was discovered that the activation of harmful instructions increases and then decreases in correlation with the "refusal direction" as the model layers deepen (they suggest that the model’s refusal of harmful instructions is mediated by a single direction) and the middle layers may play a more significant role in ensuring safety.

### v2

@@ Abstract
With the growing size of Large Vision-Language Models (LVLMs), network pruning techniques designed to compress these models for deployment in resource-constrained environments have attracted significant attention. However, we observe that pruning frequently results in a degradation in safety performance. To address this issue, we propose a novel and lightweight approach, named Hierarchical Safety Realignment (HSR). HSR operates by first quantifying the contribution of each attention head to safety, identifying the most critical ones, and then selectively restoring neurons directly within these attention heads that play a pivotal role in maintaining safety. This process hierarchically realigns the safety of pruned LVLMs, progressing from the attention head level to the neuron level. We validate HSR across various models and pruning strategies, consistently achieving notable improvements in safety performance. To the best of our knowledge, this is the first work explicitly focused on restoring safety in LVLMs post-pruning. The code will be available at https://github.com/TheShineyue/HSR.
Hierarchical Safety Realignment: Lightweight Restoration of Safety
in Pruned Large Vision-Language Models
Yue Li1thanks: Equal Contribution., Xin Yi [^1] footnotemark: 1, Dongsheng Shi1, Gerard de Melo2 Xiaoling Wang1, Linlin Wang1thanks: Corresponding Author. 1East China Normal University 2Hasso Plattner Institute/University of Potsdam {yue_li,xinyi,dongsheng}@stu.ecnu.edu.cn, gdm@demelo.org, {xlwang,llwang}@cs.ecnu.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) benefit from their massive parameter count and advanced architectures, achieving outstanding results on diverse benchmarks. Building on this success, efforts to extend LLMs into multimodal domains have made remarkable progress as well. Most current Large Vision-Language Models (LVLMs), typically composed of visual encoders, adapters, and LLM backbones (Liu et al., 2024a), have a large parameter scale and leverage image-text datasets to achieve effective multimodal alignment. To enable model deployment and application under resource-constrained environments, pruning methods (Sun et al., 2024; Frantar and Alistarh, 2023; Lee et al., 2018) compute importance scores of neurons to eliminate those deemed less important, thereby reducing the size of the model while retaining the utility to a certain extent. Such methods have seen broad adoption to boost model efficiency.
@@
Figure 1: The Average ASR (lower values indicate higher safety) of full baselines versus pruned models (50% sparsity) across safety evaluation benchmarks. The abbreviations on the x-axis, listed from left to right, correspond to the following models: Qwen2-VL, Qwen2.5-VL, LLaVA-NeXT-Mistral, LLaVA-NeXT-Vicuna, LLaVA-NeXT-Llama3, and Llama3.2-Vision.
@@
Safety is commonly defined as preventing models from following malicious instructions and generating toxic content (Bianchi et al., 2024). Recent work (Zhou et al., 2025) has revealed that certain safety heads within the attention mechanism are crucial for feature integration in safety-related tasks. Additionally, neural-level research (Wei et al., 2024) has found that certain regions within the model are critical for safety guardrails, which are separate from the utility-related regions and exhibit sparsity. A natural concern arises from the fact that, because these regions contribute minimally to utility, they are prone to being removed by pruning technologies that prioritize utility importance as a pruning metric. This removal could result in a decline in the safety of the pruned model. To verify whether this problem exists, we used the Wanda pruning method (Sun et al., 2024) to prune six mainstream LVLMs and compared their safety changes before and after pruning. The experimental results shown in Figure 1 reveal that all LVLMs exhibited varying degrees of safety degradation, with the worst-performing model showing a 15.4% safety drop and the best-performing case exhibiting a 2.8% decline. Despite the serious safety risks of pruning technologies, research on model safety restoration after pruning remains scarce.
@@
To address this problem, we propose a novel Hierarchical Safety Realignment (HSR) approach, designed to restore the safety performance degraded by pruning, without significantly increasing the pruned model’s parameter size. Our HSR apporach hierarchically realigns the safety of the pruned model from attention head level to neuron level. HSR operates in two main steps: First, we evaluate each attention head’s contribution to model safety and identify the safety-critical heads with the greatest impact. Subsequently, for these key heads, we pinpoint and restore safety-critical neurons that were pruned, effectively realigning the safety of our model.
@@
We have validated the effectiveness of our approach across various models and pruning techniques. Our proposed HSR approach successfully realigns the safety of pruned models, restoring over 27% of the lost safety in many cases and more than 14% even in the worst-case scenarios, all with lightweight modifications. Furthermore, with extensive analysis and ablation experiments, we have uncovered several key insights into model safety. These include the finding that a small subset of neurons plays a disproportionately significant role in ensuring safety, and the observation that certain neurons exist which negatively impact safety.
@@
  • We propose a novel method named Hierarchical Safety Realignment (HSR) to realign the safety of pruned LVLMs, yielding substantial safety improvements with lightweight modifications. To our knowledge, HSR is the first method specifically designed to address the safety realignment of pruned LVLMs.
@@
  • We validate the proposed method on various LVLMs using different pruning techniques. Extensive experiments demonstrate the superiority of our approach, consistently improving performance with minimal neuron restoration.
@@ 2.2.1 Quantifying Neuron Importance
    This equation is the first-order Taylor approximation to the change of the loss when the weight entry $\mathbf{W}_{ij}$ is set to zero.
    Following the experimental setup described by Wei et al. (2024), for a given calibration dataset $D$, we use
@@ 3 Experimental Setup
During the evaluation phase, we employ the following benchmarks: (1) For utility assessment, MMbench (Liu et al., 2025) and DocVQA (Mathew et al., 2021) are utilized; (2) For safety evaluation, safebench-mini (Ying et al., 2024a) and the harmful subset of $\text{C}h^{3}\text{Ef}$ (Shi et al., 2024) are adopted. To ensure fair comparison and reproducibility, all evaluations are conducted under strict zero-shot settings with greedy decoding strategies.
@@
Our experiments involved six mainstream LVLMs, including three variants of LLaVA-NeXT [^1] https://llava-vl.github.io/blog/2024-05-10-llava-next-stronger-llms/, built on different language models: Vicuna, Mistral, and Llama3. Additionally, we evaluated Qwen2.5-VL [^2] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct, Qwen2-VL (Wang et al., 2024a), and Llama-3.2-Vision [^3] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision-Instruct. All models have parameter counts ranging from 7B to 11B.
@@
Three evaluation metrics we use are Attack Success Rate (ASR), Average Normalized Levenshtein Similarity (ANLS) and Accuracy (Acc). ASR is used to evaluate the safety of the model. The smaller the ASR, the better the safety. We use Llama-Guard-3-Vision [^4] https://huggingface.co/meta-llama/Llama-Guard-3-11B-Vision to determine whether the response is safe. We use ANLS and Acc as the evaluation indicators of DocVQA and MMbench respectively. Following Mathew et al.’s (2021) proposal, ANLS can ensure that minor answer mismatches stemming from OCR errors are not severely penalized. Additionally, in order to more intuitively and fairly reflect the performance improvement of our method, we provide the Ratio of model Safety Realignment (RSR) as follows:
@@
to quantify the ratio of restored safety to lost safety. More relevant details are reported in Appendix  A.
@@ 4 Experimental Results
Regarding utility, both Qwen2-VL and the aforementioned Qwen2.5-VL show improved performance, while other models experience slight declines. This difference may be attributed to the Qwen series employing grouped query attention, which features the largest number of queries per group and the fewest heads (as shown in Appendix B), making each neuron’s contribution (both utility and safety) particularly significant.
Regarding the neuron restoration ratio, for LLaVA-NeXT-Vicuna, it is significantly higher than for others. This may be because other models are based on the group attention mechanism, and the safety realignment brought by the restoration can affect a wider range.
@@
Moreover, the utility of realigned models exhibits a slight reduction at 40% sparsity while showing progressive improvement at 50%~60% sparsity levels, peaking at 60% sparsity. This phenomenon may stem from safety-critical neurons possessing greater safety significance than utility value, whereas in heavily pruned models (50% ~60% sparsity) with substantial utility degradation, the utility of the safety-critical neurons can still bring about some improvement.
@@
Next, we analyze the impact of the hyper-parameter $\text{p}_{\max}$ as shown in Table 6. The results reveal that utility initially increases and then decreases, reaching its peak at 0.55. This underscores the importance of carefully designing $\text{p}_{\max}$ to regulate safety-related neurons. Regarding the trend of safety first deteriorating and then improving, this suggests that some neurons contribute significantly to both safety and utility, while others contribute minimally to both. By selecting an appropriate $\text{p}_{\max}$, we can effectively exclude the latter group.
@@
Finally, we examine the impact of the hyper-parameter $h$ as presented in Table 7. Since each group attention head in LLaVA-NeXT-Mistral corresponds to 4 query matrices, experiments are conducted in multiples of 4. The results show that safety performs best when $h$ equals 4, followed by significant fluctuations. Utility also shows considerable variability. This may be due to the uneven distribution of neurons across head, with some contributing significantly to both safety and utility, and others contributing little. This highlights the need for more fine-grained control to address this variability effectively.
@@
  • Lightweight Implementation: HSR can achieve safety realignment of pruned model while preserving the current sparsity of the pruned model as much as possible. For Qwen2.5-VL and LLaVA-NeXT-Llama3, the scale of neurons restored by HSR-a is 18 and 650 times that of HSR, respectively.
  • Enhanced Effectiveness: Certain neurons have been found to adversely affect safety. By filtering them out through set operations, we ensure robust safety realignment efficacy. Directly repairing the entire head makes the safety worse (ASR increased > 0.2). This also shows again that there may be some neurons that are harmful to safety.
@@ 5 Further Analysis
We analyze the total Ships of each model and the extent of safety degradation after pruning, finding a strong positive correlation between them. Spearman’s rank correlation coefficient $\rho$ (Spearman, 1904) is a nonparametric statistical test measuring the correlation between the ranks of two variables:
@@ 6 Related Work
Many studies have researched methods to compromise the safety of LVLMs. Gong et al. (2025) introduced FigStep, which converts prohibited content into images using typesetting to bypass safety alignment. Ying et al. (2024b) proposed BAP, a jailbreak attack method that jointly optimizes text and visual prompts. In response, research on defending against such attacks and enhancing model safety has also emerged. Liu et al. (2024b) improved defense against harmful images by incorporating a security module via a two-stage training process. Meanwhile, Wang et al. (2024b) proposed AdaShield, which protects Multimodal Large Language Models from structure-based jailbreak attacks by adding a defense hint to the input, without requiring model fine-tuning or additional module training.
@@
As model sizes continue to grow, pruning techniques (Sung et al., 2024; Cao et al., 2024) for compressing neural networks by removing neurons have attracted significant attention. These techniques can be broadly categorized into structured and unstructured pruning. Structured pruning (Ashkboos et al., 2024; An et al., 2024) has the advantage of accelerating pruned models on standard hardware without relying on specialized support (Zhu et al., 2024), while unstructured pruning (Lee et al., 2019; Sun et al., 2024; Frantar and Alistarh, 2023) helps preserve performance at higher sparsity levels. For pruned models, Jin et al. (2022) observed that pruning introduces additional regularization, reducing accuracy loss on noisy examples in dense models. Hasan et al. (2024) noted improved model safety at low sparsity, attributing it to sharper attention. However, the safety degradation caused by pruning at slightly higher sparsity has been overlooked, motivating our research on methods to realign the safety of pruned models.
@@ 7 Conclusion
In this paper, we propose a novel Hierarchical Safety Realignment (HSR) approach to mitigate the overemphasis on neuron utility in pruning methods, which may lead to a significant degradation in the safety of pruned models. Specifically, HSR first identifies the safety-critical attention heads that contribute significantly to safety at the attention head level, and subsequently restores the safety-critical neurons that were pruned within those attention heads. Extensive experiments on multiple mainstream LVLMs and pruning methods demonstrate that HSR achieves lightweight yet effective safety realignment by leveraging the fact that only a relatively small number of neurons significantly contribute to model safety. We hope that our safety realignment approach can facilitate the deployment of compact and reliable models.
@@ Appendix E Analysis and visualization of the overlap between $S^{s}(\text{q})$ and $S^{u}(\text{p})$
  • For Q, K, V and O matrices, the degree of differentiation between safety and utility increases initially and then decreases as the layers deepen. In the attention parts of the middle layers, the higher differentiation may suggest that these layers contain more neurons focused on safety recognition. This observation is similar to the findings from prior work (Arditi et al., 2024), where it was discovered that the activation of harmful instructions increases and then decreases in correlation with the "refusal direction" as the model layers deepen (they suggest that the model’s refusal of harmful instructions is mediated by a single direction) and the middle layers may play a more significant role in ensuring safety.